# Data $\times$ LLM: From Principles to Practices

### v1

@@ 1.1 Techniques of DATA4LLM
$\bullet$ LLM Model Data Fault Tolerance. LLM model data fault tolerance aims to enhance the ability to recover from system failures during model training. Relevant techniques include (1) checkpointing [292, 195, 410, 395], which stores checkpoints across a hierarchical storage system; and (2) redundant computation [387, 187, 147], which replicates the state data of LLMs to support rapid fault recovery.
@@ 2.2.2 Data Deduplication
$\bullet$ DotHash [299]. Moreover, to further improve the deduplication accuracy and efficiency, DotHash [299] assumes that uniformly sampled vectors in high-dimensional space are quasi-orthogonal. It encodes sample elements into fixed-length basis vectors, and treats the samples as combinations of these basis vectors, thus the dot product of these vectors is an unbiased estimate of their intersection. For example, given the set $a=\sum_{a\in A}\psi(a)\quad$ and the set $\quad b=\sum_{b\in B}\psi(b)$, the intersection is calculated by $\mathbb{E}[a\cdot b]=|A\cap B|$.
However, [121] found that DotHash would perform badly if the basis vector has a shorter length than the number of basis vectors, where quasi-orthogonal no longer holds.
Approximate Frequency-based Down-Weighting. To prevent the loss of potentially valuable information by retaining only one sample and removing the rest, SoftDeDup [168] deduplicates by reweighting samples, where samples with higher commonness will be assigned higher sampling weights, and others with lower commonness will be assigned lower sampling weights. Specifically, SoftDeDup computes the frequency of each n-gram across all the samples, and then calculates the commonness of each sample by multiplying the frequencies of all the n-grams that appear in the document. The samples with higher commonness would be seen as more likely to be duplicate and thus be down-weighted.
@@
SemDeDup [46] identifies semantic duplicates by clustering embeddings and deduplicating those with high cosine similarities. It first encodes each data point into an embedding by leveraging the OPT [469] text encoder and the CLIP [326, 183] image encoder, and clusters the embeddings with k-means, so one can save time by finding duplicates within the cluster rather than the whole vector space. Then, within each cluster, it searches for semantic duplicates with cosine similarity above the pre-defined threshold. Finally, within each group of duplicates, it retains the example with the lowest cosine similarity to the cluster centroid and removes the rest. As a multi-modal method, it can be applied to both text and image data, making it possible to deduplicate image data. In practice, [45] leverages SemDeDup to deduplicate the image-text pair dataset LAION-400M [344].
Like MinHash, SemDeDup also has many variants for performance improvement. [390] combines SemDeDup with the Self-Supervised Learning (SSL) Prototypes metric, such as representing unlabeled data in terms of their cluster centroids in an embedding space, for data pruning to further remove semantic duplicates that cannot be removed by MinHash. FairDeDup [364] modifies the logic of SemDeDup to improve the representation of underrepresented sensitive groups by prioritizing the retention of data samples that align with sensitive concepts defined through user-provided prototypes, such as demographic subgroups. Within each cluster, instead of selecting the farthest sample from the centroid, it selects the sample that maximizes similarity to the least-represented group in the cluster to prevent samples with sensitive concepts from being pruned.
Non-Text Data Deduplication. As LLMs are increasingly applied to multimodal tasks (e.g., image-text retrieval, visual question answering), non-text data types such as images are becoming integral to LLM training datasets, necessitating dedicated deduplication techniques. Like texts, images can be encoded into embeddings through neural networks designed for images, like CNN, after which embedding-based deduplication methods can be applied. SemDedup [46] adopts a semantic-based method by computing cosine similarity between image embeddings; two images are considered duplicates if their similarity exceeds a predefined threshold, which is tuned to balance detection precision and recall. In contrast, MINT-1T employs a hash-based approach, using SHA256 checksums to identify and remove exact duplicates efficiently. Meanwhile, the DataComp pipeline [146] leverages the CNN-based near-duplicate detector [453] to eliminate subtle duplicates and prevent evaluation set leakage. Models trained on these deduplicated image sets exhibit improved performance over baselines such as CLIP [326] for higher precision and recall.
@@ 2.2.4 Data Selection
$\bullet$ Bayes-based Selection: CoLoR-filter [80] formulates pre-training subset selection as a Bayesian optimization problem, which selects a subset $S$ by minimizing downstream loss or maximizing downstream task performance $\text{Pr}(D_{\text{down}}|S)$. It uses two auxiliary models: A “prior” model ($\theta_{\text{prior}}$) trained on a large general dataset $D_{\text{down}}$ and a “conditional” model ($\theta_{\text{prior}}$) fine-tuned on the union of the large general dataset and a small downstream dataset $D_{\text{prior+down}}$. The selection criterion for a data point $x_{i}$ is the conditional loss reduction (CoLoR): $\text{CoLoR}(x_{i})=-\log\Pr(x_{i}|\theta_{\text{prior+down}})-(-\log\Pr(x_{i}|\theta_{\text{prior}}))$. The key idea is to score data points based on the difference in their likelihood under these two models, selecting those that are more likely under the conditional model. This conditional loss reduction (CoLoR) criterion helps identify data that is particularly useful for the target tasks.
@@
$\bullet$ Linear Search. Model-Aware Dataset Selection with Datamodels (DsDm) [130] selects the optimal subset of training data that minimizes the model’s loss on target tasks by employing linear datamodel [185], a parameterized function that maps a subset of training data to the model outputs for the specified target, to estimate how the inclusion of each training example would affect the model’s loss on the target, avoiding infeasible computations. In practice, a linear datamodel $\tau_{\theta_{x}}(1_{S})=\theta_{x}^{\top}1_{S}$ with parameters $\theta_{x}$ and a characteristic vector $1_{S}$ (a binary vector indicating which data points are in $S$) is adopted to map the subset $S$ to the model loss on a sample $x$ through $L_{x}(S)=\mathbb{E}[\ell(x;A(S))]$. For each target, the characteristic vector $1_{S}$ is adjusted to reflect the subset, and the parameters $\theta_{x}$ are estimated using a regression loss function like mean squared error over the training subset. After training, the datamodel selects the subset $S$ of the size $k$ that minimizes the loss $\hat{L}_{D_{\text{targ}}}(S)=\frac{1}{n}\sum_{i=1}^{n}\tau_{\theta_{x_{i}}}(1_{S})$ for the target task.
@@
$\bullet$ Gradient-Influence Search. Low-rank Gradient Similarity Search (LESS) [424] identifies the most impactful subset of data for fine-tuning LLMs by analyzing gradient similarities. It first fine-tunes the model on a random subset (e.g., 5% of data) for a few epochs using LoRA to reduce trainable parameters and accelerate gradient computation, and saves the checkpoints after each epoch. Next, LESS computes Adam LoRA gradients for each training sample, projects them into lower-dimensional gradient features via random projection, and stores them in a gradient datastore. For downstream tasks, it calculates gradient features of few-shot validation samples and estimates the influence of each training sample $\bm{z}$ on a validation sample $\bm{z}^{\prime}$ using cosine similarity: $\text{Inf}_{\text{Adam}}(\bm{z},\bm{z}^{\prime})\triangleq\sum_{i=1}^{N}\bar{\eta}_{i}\cos(\nabla\ell(\bm{z}^{\prime};\bm{\theta}_{i}),\Gamma(\bm{z},\bm{\theta}_{i}))$, where $\Gamma(\bm{z},\bm{\theta})$ is the Adam update. The training samples with the highest influence scores are selected for fine-tuning. Experiments show that models trained on this subset outperform those trained on full data, and the approach can scale to larger models.
@@
$\bullet$ Kernel-Density Regularization. Task-Specific Data Selection (TSDS) [270] identifies high-quality pre-training or fine-tuning data for particular tasks by balancing two objectives: $(i)$ distribution alignment with the target task data and $(ii)$ diversity to avoid near-duplicates. It accomplishes this via an optimal transport formulation enhanced by kernel density estimation (KDE) regularization. Concretely, one begins with a small set of query examples $Q=\{q_{i}\}_{i=1}^{M}$ (representing the target task) and a large candidate pool $D=\{x_{j}\}_{j=1}^{N}$. Both $Q$ and $D$ are then embedded into a shared metric space (e.g., using gradient-based or semantic embeddings). The optimization for distribution alignment is conducted by solving for probability mass $\gamma_{ij}$ (transported from $q_{i}$ to $x_{j}$): $\min_{\gamma\in\mathbb{R}_{\geq 0}^{M\times N}}\frac{\alpha}{C}\sum_{i=1}^{M}\sum_{j=1}^{N}\gamma_{ij}d_{ij}+(1-\alpha)G_{\text{KDE}}(\gamma)\quad\text{s.t.}\quad\sum_{j=1}^{N}\gamma_{ij}=\frac{1}{M},\forall i\in[M]$, where $d_{ij}$ is the distance between $q_{i}$ and $x_{j}$ in the metric space, and $G_{\text{KDE}}(\gamma)$ is the regularization term that adds diversity and penalizes over-density using KDE estimates: $G_{\text{KDE}}(\gamma)=M\max_{i,j}\rho_{j}\left|\gamma_{ij}-\frac{1/\rho_{j}}{M\sum_{j^{\prime}}1/\rho_{j^{\prime}}}\right|$, where $\rho_{j}=\sum_{x^{\prime}\in D}(1-f(x_{j},x^{\prime})^{2}/h^{2}$ is the density estimate for candidate $x_{j}$ (higher for near-duplicates). Afterwards, it selects samples by sampling $x_{j}$ with probability $p_{j}=\sum_{i}\gamma_{ij}^{*}$.
@@ 2.2.5 Data Mixing
TABLE VI: Comparison of Data Mixing Methods for LLM
@@
Heuristic Optimization Pre-training Multi-source Data Adjusting Intuitive and easy to implement, suitable for rapid experimentation
@@
Entropy-Based Mixing [153] Low computation cost with quality quantification by entropy
Bilevel Optimization Pre-training Calculate domain contribution by Requires a proxy model, performances well in OOD datasets
@@
Fine-tuning Task-level DRO [279] Performs well with problems caused by significant
different loss across datasets
Model-Based Pre-training Linear Regression Model [264] Only 10% of DoReMi’s [428] computational resources are required.
Supports simultaneous training hundreds of small models to accelerate the optimization
Pre-training Bivariate Data Mixing Law [153] Avoid iterative training of proxy models, resulting in lower
computational costs. Shows relation between loss and training steps additionally
Continual Pre-training Chinchilla Scaling Law [324] Supports knowledge transfering to new domains, reducing training costs by over 95%
Pre-training Exponential Functions [447] Supports datasets without explicit domain division.
Continual Pre-training Power-law Function [161] Compared to [324] which is single-objective optimization
essentially, [161] ensures that domain performance improvement
@@
Pre-training Classification Model [252] Reverse engineering for finding the data recipe of LLMs
@@ 2.2.6 Data Distillation and Synthesis
$\bullet$ Task-Specific Prompt Distillation. To significantly reduce inference costs and latency while maintaining performance, [357] employs task-specific prompts: (1) Chain-of-Density (CoD): Iteratively adds entities to summarize for enhanced density. (2) Chain-of-Thought (CoT): Guides reasoning tasks (e.g., math) through stepwise logic. Using GSM8K [107] data and Llama-3.1-405B-Instruct, synthetic data is generated for fine-tuning smaller models (Llama-3.1-8B/70B-Instruct) paired with simplified prompts, balancing efficiency and task specialization.
@@ 2.2.7.1 Typical data processing frameworks
(1) Data-juicer [90] is an open-source data processing framework with customization, high-quality, diverse, and efficient data recipes. It has 50 built-in operators for tasks like formatting, mapping, filtering, and deduplication. Besides, Data-juicer incorporates visualization and auto-evaluation capabilities to provide timely feedback, and is optimized for distributed computing to handle large-scale data efficiently.
(2) Dataverse [306] is an open-source pipeline to streamline the development of custom ETL (Extract-Transform-Load) pipelines using a block-based interface that allows users to intuitively customize their ETL pipelines by simply adding, removing, or reshuffling blocks. Dataverse includes a wide range of native operations for data processing, such as deduplication, decontamination, bias mitigation, and toxicity reduction, and supports multi-source data ingestion. Similar to Data-juicer, Dataverse integrates with Apache Spark for distributed processing and supports AWS integration for cloud scalability.
(3) [373] introduces a data processing framework that allows users to customize data processing pipelines without manual coding. It has two main modules: (1) The processing module consists of operators for data reformatting (read and import strctured data), cleaning (removed undesired data such as HTML tags and translate text), filtering, and deduplication (using MinHashLSH in Section 2.2.2); (2) The analyzing module supports probing and evaluation of the refined data.
@@ 2.2.7.3 Orchestration of data pipelines
The above data pipelines are mostly designed by experience. Instead, Data-Juicer Sandbox [91] proposes a “Probe-Analyze-Refine” workflow, which involves systematically exploring the impact of various data processing operations and their orders on model performance, combining effective operations into data recipes, and optimizing data utilization through a dual analysis focusing on duplication and diversity. The orchestrated pipelines are validated through applications on state-of-the-art models like Mini-Gemini (for image-to-text generation) and EasyAnimate (for text-to-video generation).

### v2

@@ 1.1 Techniques of DATA4LLM
$\bullet$ LLM Model Data Fault Tolerance. LLM model data fault tolerance aims to enhance the ability to recover from system failures during model training. Relevant techniques include (1) checkpointing [293, 196, 411, 396], which stores checkpoints across a hierarchical storage system; and (2) redundant computation, which leverages redundant states of LLM in parallel training (e.g., pipeline parallelism [388], hybrid parallelism [188, 148]) to support rapid fault recovery.
@@ 2.2.2 Data Deduplication
$\bullet$ DotHash [300]. Moreover, to further improve the deduplication accuracy and efficiency, DotHash [300] assumes that uniformly sampled vectors in high-dimensional space are quasi-orthogonal. It encodes each sample into a combination of sample elements represented as fixed-length basis vectors, and the dot product of these vectors is an unbiased estimate of their intersection. For example, given two samples with their element basis vectors $a=\sum_{a\in A}\psi(a)\quad$ and $\quad b=\sum_{b\in B}\psi(b)$, the intersection is calculated by $\mathbb{E}[a\cdot b]=|A\cap B|$.
However, [122] found that DotHash performs badly if the length of the basis vector is lower than the number of basis vectors, where quasi-orthogonal no longer holds.
Approximate Frequency-based Down-Weighting. To prevent the loss of potentially valuable information by retaining only one sample and removing the rest, SoftDeDup [169] deduplicates by reweighting samples, where samples with higher commonness are assigned lower sampling weights. Specifically, SoftDeDup computes the frequency of each n-gram across all the samples and calculates the commonness of each sample by multiplying the frequencies of all the n-grams that appear in the document. Samples with higher commonness are more likely to be duplicates and thus be down-weighted.
@@
SemDeDup [46] identifies semantic duplicates by clustering embeddings and deduplicating those with high cosine similarities. It first encodes each sample into an embedding by leveraging the OPT [471] text encoder and the CLIP [327, 184] image encoder, and clusters the embeddings with K-means, so one can save time by finding duplicates within the cluster rather than the whole vector space. Then, within each cluster, it searches for semantic duplicates with cosine similarity above the pre-defined threshold. Finally, within each group of duplicates, it retains only the sample closest to the cluster centroid. As a multi-modal method, it can be applied to both text and image data, making it possible to deduplicate image data. In practice, [45] leverages SemDeDup to deduplicate the image-text pair dataset LAION-400M [345].
Like MinHash, SemDeDup also has many variants for performance improvement. [391] combines SemDeDup with the Self-Supervised Learning (SSL) Prototypes metric, which clusters the samples and retains the samples in each cluster based on their distance to their corresponding cluster centroid, where the samples closer to the centroid are more likely to be removed. FairDeDup [365] modifies the logic of SemDeDup to improve the representation of underrepresented sensitive groups by prioritizing the retention of samples that align with sensitive concepts defined through user-provided prototypes, such as demographic subgroups. Within each cluster, instead of selecting the farthest sample from the centroid, it selects the sample that maximizes similarity to the least-represented group in the cluster to prevent samples with sensitive concepts from being pruned.
Non-Text Data Deduplication. As LLMs are increasingly applied to multimodal tasks (e.g., image-text retrieval, visual question answering), non-text data types such as images are becoming integral to LLM training datasets, necessitating dedicated deduplication techniques. Similar to texts, images can also be encoded into embeddings through neural networks designed for image-like data such as CNN, after which embedding-based deduplication methods can be applied. SemDedup [46] adopts a semantic-based method by computing cosine similarity between image embeddings; two images are considered duplicates if their similarity exceeds a predefined threshold, which is tuned to balance detection precision and recall. In contrast, MINT-1T employs a hash-based approach, using SHA256 checksums to identify and remove exact duplicates efficiently. Meanwhile, the DataComp pipeline [147] leverages the CNN-based near-duplicate detector [454] to eliminate subtle duplicates and prevent evaluation set leakage. Models trained on these deduplicated image sets exhibit improved performance over baselines such as CLIP [327] for higher precision and recall.
@@ 2.2.4 Data Selection
$\bullet$ Bayes-based Selection: CoLoR-filter [80] formulates pre-training subset selection as a Bayesian optimization problem, which selects a subset $S$ by maximizing downstream likelihood $\text{Pr}(D_{\text{down}}|S)$. It uses two auxiliary models: A “prior” model ($\theta_{\text{prior}}$) trained on a large general dataset $D_{\text{down}}$ and a “conditional” model ($\theta_{\text{prior}}$) fine-tuned on the union of the large general dataset and a small downstream dataset $D_{\text{prior+down}}$. The selection criterion for a data point $x_{i}$ is the conditional loss reduction (CoLoR): $\text{CoLoR}(x_{i})=-\log\Pr(x_{i}|\theta_{\text{prior+down}})-(-\log\Pr(x_{i}|\theta_{\text{prior}}))$. The key idea is to score samples based on the likelihood difference between these two models and select the ones that exhibit higher likelihood under the conditional model and larger conditional loss reduction.
@@
$\bullet$ Linear Search. Model-Aware Dataset Selection with Datamodels (DsDm) [131] selects the optimal subset of training data that minimizes the model’s loss on target tasks by employing linear datamodel [186], a parameterized function that maps a subset of training data to the model outputs for the specified target, to estimate how the inclusion of each training sample would affect the model’s loss on the target, reducing computational overhead. In practice, a linear datamodel $\tau_{\theta_{x}}(1_{S})=\theta_{x}^{\top}1_{S}$ with parameters $\theta_{x}$ and a characteristic vector $1_{S}$ (a binary vector indicating which samples are in $S$) is adopted to map the subset $S$ to the model loss on a sample $x$ through $L_{x}(S)=\mathbb{E}[\ell(x;A(S))]$. For each target, the characteristic vector $1_{S}$ is adjusted to reflect the subset, and the parameters $\theta_{x}$ are estimated using a regression loss function like mean squared error over the training subset. After training, the datamodel selects the subset $S$ of the size $k$ that minimizes the loss $\hat{L}_{D_{\text{targ}}}(S)=\frac{1}{n}\sum_{i=1}^{n}\tau_{\theta_{x_{i}}}(1_{S})$ for the target task.
@@
$\bullet$ Gradient-Influence Search. Low-rank Gradient Similarity Search (LESS) [425] identifies the most impactful subset of data for fine-tuning LLMs by analyzing gradient similarities. It first fine-tunes the model on a random subset (e.g., 5% of data) for a few epochs using LoRA to reduce trainable parameters and accelerate gradient computation, and saves the checkpoints after each epoch. Next, LESS computes Adam LoRA gradients for each training sample, projects them into lower-dimensional gradient features via random projection, and stores them in a gradient datastore. For downstream tasks, it calculates gradient features of few-shot validation samples and estimates the influence of each training sample $\bm{z}$ on a validation sample $\bm{z}^{\prime}$ using cosine similarity: $\text{Inf}_{\text{Adam}}(\bm{z},\bm{z}^{\prime})\triangleq\sum_{i=1}^{N}\bar{\eta}_{i}\cos(\nabla\ell(\bm{z}^{\prime};\bm{\theta}_{i}),\Gamma(\bm{z},\bm{\theta}_{i}))$, where $\Gamma(\bm{z},\bm{\theta})$ is the Adam update. The training samples with the highest influence scores are selected for fine-tuning.
@@
$\bullet$ Kernel-Density Regularization. Task-Specific Data Selection (TSDS) [271] identifies high-quality pre-training or fine-tuning data for particular tasks by balancing two objectives: $(i)$ distribution alignment with the target task data and $(ii)$ diversity to avoid near-duplicates, accomplished via kernel density estimation (KDE) regularization. Concretely, one begins with a small set of target task samples $Q=\{q_{i}\}_{i=1}^{M}$ and a large candidate pool $D=\{x_{j}\}_{j=1}^{N}$, both of which are embedded into a shared metric space (e.g., using gradient-based or semantic embeddings). The optimization for distribution alignment is conducted by solving for probability mass $\gamma_{ij}$ (transported from $q_{i}$ to $x_{j}$): $\min_{\gamma\in\mathbb{R}_{\geq 0}^{M\times N}}\frac{\alpha}{C}\sum_{i=1}^{M}\sum_{j=1}^{N}\gamma_{ij}d_{ij}+(1-\alpha)G_{\text{KDE}}(\gamma)\quad\text{s.t.}\quad\sum_{j=1}^{N}\gamma_{ij}=\frac{1}{M},\forall i\in[M]$, where $d_{ij}$ is the distance between $q_{i}$ and $x_{j}$ in the metric space, and $G_{\text{KDE}}(\gamma)$ is the regularization term that adds diversity and penalizes over-density using KDE: $G_{\text{KDE}}(\gamma)=M\max_{i,j}\rho_{j}\left|\gamma_{ij}-\frac{1/\rho_{j}}{M\sum_{j^{\prime}}1/\rho_{j^{\prime}}}\right|$, where $\rho_{j}=\sum_{x^{\prime}\in D}(1-f(x_{j},x^{\prime})^{2}/h^{2}$ is the density estimate for candidate $x_{j}$ (higher for near-duplicates). Afterwards, it samples $x_{j}$ with probability $p_{j}=\sum_{i}\gamma_{ij}^{*}$.
@@ 2.2.5 Data Mixing
TABLE VI: Comparison of Data Mixing Methods for LLM.
@@
Heuristic Optimization Pre-training Multi-Source Data Adjusting Intuitive and easy to implement, suitable for rapid experimentation.
@@
Entropy-Based Mixing [154] Low computation cost with quality quantification by entropy.
Bilevel Optimization Pre-training Calculate domain contribution by Requires a proxy model, performances well in OOD datasets.
@@
Fine-tuning Task-level DRO [280] Perform well with problems with significant
loss difference across datasets
Model-Based Optimization Pre-training Linear Regression Model [265] Only 10% of DoReMi’s [429] computational resources are required.
Simultaneously train hundreds of small models to accelerate optimization.
Pre-training Bivariate Data Mixing Law [154] Avoid iterative training of proxy models (low computational costs).
Show relation between loss and training steps.
Continual Pre-training Chinchilla Scaling Law [325] Support knowledge transferring to new domains ($\downarrow$ over 95% training costs).
Pre-training Exponential Functions [448] Support datasets without explicit domain division.
Continual Pre-training Power-law Function [162] Compared to single-objective optimization like [325]
[162] ensures that domain performance improvement
@@
Pre-training Classification Model [253] Reverse engineering for finding the suitable data recipe of LLMs.
@@ 2.2.6 Data Distillation and Synthesis
$\bullet$ Task-Specific Prompt Distillation. To significantly reduce inference costs and latency while maintaining performance, [shirgaonkar202x4knowledge] employs task-specific prompts: (1) Chain-of-Density (CoD): Iteratively adds entities to summarize for enhanced density. (2) Chain-of-Thought (CoT): Guides reasoning tasks (e.g., math) through stepwise logic. Using GSM8K [107] data and Llama-3.1-405B-Instruct, synthetic data is generated for fine-tuning smaller models (Llama-3.1-8B/70B-Instruct) paired with simplified prompts, balancing efficiency and task specialization.
@@ 2.2.7.1 Typical data processing frameworks
(1) Data-juicer [90] is an open-source framework designed for customizable, high-quality, and efficient data processing. It offers a diverse range of pre-built data processing operators such as data formatting, mapping, filtering, and deduplication. Additionally, the framework features visualization and automatic evaluation, enabling users to receive immediate feedback on their data pipeline. To manage large-scale datasets effectively, Data-juicer is optimized for distributed computing, ensuring robust performance and scalability.
(2) Dataverse [307] is an open-source framework designed to simplify custom ETL (Extract-Transform-Load) pipeline development through an easy-to-use block-based interface that enables users to easily customize by adding, removing, or rearranging blocks. The platform offers a diverse range of pre-built data processing operators, including deduplication, decontamination, bias mitigation, and toxicity reduction, while also supporting the integration of data from multiple sources. Similar to Data-juicer, Dataverse integrates with Apache Spark for distributed processing and supports AWS integration for cloud scalability.
(3) [374] introduces a data processing framework that allows users to customize data processing pipelines using a comprehensive suite of operators categorized in two main modules: (1) The processing module consisting of data reformatting (read and import strctured data), cleaning (removed undesired data such as HTML tags and translate text), filtering, and deduplication (using MinHashLSH in Section 2.2.2) operators; (2) The analyzing module featuring refined data probing and automatic evaluation.
@@ 2.2.7.3 Orchestration of data pipelines
The above data pipelines are mostly designed by experience. Instead, Data-Juicer Sandbox [91] proposes a “Probe-Analyze-Refine” workflow, which involves systematically exploring the impact of various data processing operations and their orders on model performance, combining effective operations into data recipes, and optimizing data utilization through duplication analysis and diversity analysis. The orchestrated pipelines are validated through applications on state-of-the-art models like Mini-Gemini (for image-to-text generation) and EasyAnimate (for text-to-video generation).